OpenStudy (anonymous):

factor using the GCF: 8x^2y-8xy

OpenStudy (callisto):

\[8x^2y-8xy = 8(x)(x)(y) - 8(x)(y)\] What are the common factors of the two terms?

OpenStudy (anonymous):

8x^2y

OpenStudy (callisto):

Nope... 8 (x) (x) (y) | 8 (x)(y) One of the common factors is 8. There are two more, what are they?

OpenStudy (anonymous):

2 and 4

OpenStudy (callisto):

I meant for the terms 8x^y and 8xy..

OpenStudy (anonymous):

x and y

OpenStudy (callisto):

Yes. So, the GCF is 8xy, agree?

OpenStudy (anonymous):

yup

OpenStudy (callisto):

Now, take out the common factors and group the rest 8x^2y - 8xy = 8xy ( ___ - 1) Can you fill in the blanks?

OpenStudy (anonymous):

8xy(x-1)

OpenStudy (callisto):

Yes! :)

OpenStudy (anonymous):

so thats the answer??

OpenStudy (callisto):

Yes!

OpenStudy (anonymous):

ok thx i gtg cya

OpenStudy (callisto):

Welcome~ Bye for now!